### v1

@@ Abstract
I suggest the double-degenerate (DD) scenario with a merger-to-explosion delay (MED) time (the DD-MED scenario) of about 1-2 years to explain the rare properties of the recently analyzed type Ia supernova (SN Ia) SN 2020aeuh. The rare properties are the SN Ia ejecta interacting with a carbon-oxygen (CO)-rich circumstellar material (CSM) at approximately 50 days post-explosion. In this DD-MED scenario, two massive CO white dwarfs (WDs), masses of $M_{1}\simeq 1.1M_{\odot}$ and $M_{2}\simeq 1M_{\odot}$, merge to leave a rapidly rotating lonely WD of about the Chandrasekhar mass. The merger process ejects $M_{\rm CSM}\simeq 0.7M_{\odot}$ to form a nonspherical CO-rich CSM. At the explosion, there is a lonely WD and a detached hydrogen- and helium-deficient CSM. Studies proposed the other lonely WD scenario, the core-degenerate (CD) scenario, to explain several specific SNe Ia and SN Ia remnants. SN 2020aeuh is the first particular SN Ia that is attributed to the DD-MED scenario. Besides being slightly brighter than typical SNe Ia and the CSM interaction, SN 2020aeuh is a normal SN Ia. Therefore, this study strengthens the claim that the lonely WD scenarios, i.e., the DD-MED and CD scenarios, account for most, if not all, normal SNe Ia; all SN Ia scenarios, whether lonely WD or not, might contribute to peculiar SNe Ia.
@@ 1 INTRODUCTION
The old classification of type Ia supernovae (SNe Ia) scenarios, into the double-degenerate (DD) and single-degenerate (SD) scenarios, is outdated (despite many papers still mentioning this classification as if it is applicable). There are more scenarios that any modeling of an SN Ia must consider. However, there is no consensus on the classification of these scenarios as reviews show (Maoz et al. 2014; Maeda, & Terada 2016; Hoeflich 2017; Livio & Mazzali 2018; Soker 2018, 2019a, 2024a; Wang 2018; Jha et al. 2019; Ruiz-Lapuente 2019; Ruiter 2020; Aleo et al. 2023; Liu, Röpke, & Han 2023; Vinkó, Szalai, & Könyves-Tóth 2023; Ruiter & Seitenzahl 2025). No scenario is entirely free of difficulties in accounting for observations, with some encountering a few challenges, while others face many challenges. Some observations challenge most (e.g., Pearson et al. 2024) or all (e.g., Wang et al. 2024) scenarios. A challenging observation, for example, is the relation of the $\gamma$-ray escape time to the 56Ni mass (e.g., Schinasi-Lemberg & Kushnir 2025; Sharon & Kushnir 2025; Sharon, Kushnir, & Schinasi-Lemberg 2025). Studies have added more sub-scenarios or branches of the main scenarios, including the HeCO hybrid channel of the DD scenario (e.g., Perets et al. 2019; Zenati, Toonen, & Perets 2019; Zenati et al. 2023 ) common-envelope wind model (Meng & Podsiadlowski 2017; Cui et al. 2022; Wang & Meng 2025), the carbon–oxygen–neon white dwarf (WD) with a main sequence companion channel of the SD scenario (e.g., Guo et al. 2025), and new mass transfer prescriptions (e.g., Li, Liu, & Wang 2023). Studies have examined most of these scenarios in recent years, with no emerging leading scenario(s) in the consensus (e.g., Bora et al. 2024; Boos, Townsley, & Shen 2024; Bregman et al. 2024; Casabona & Fisher 2024; DerKacy et al. 2024; Joshi, Strolger, & Zenati 2024; Ko et al. 2024; Kobashi et al. 2024; Lim et al. 2024; Mehta et al. 2024; Palicio et al. 2024; Phillips et al. 2024; Soker 2024b; Uchida et al. 2024; Burmester et al. 2025; Gaba et al. 2025; Glanz et al. 2025; Griffith et al. 2025; Ito et al. 2025; Iwata & Maeda 2025; Magee et al. 2025; O’Hora et al. 2025; Pan, Ruiz-Lapuente, & González Hernández 2025; Pollin et al. 2025; Simotas, Bildsten, & Prust 2025; Wang, Chen, & Pan 2025, for a limited list of papers since 2024).
@@ 2 The DD-MED scenario for SN 2020aeuh
Tsalapatas et al. (2025) estimate the CSM shell to be from $R_{\rm CSM,inner}\simeq 3\times 10^{15}{~{}\rm cm}$ to $R_{\rm CSM,outer}\simeq 9\times 10^{15}{~{}\rm cm}$. For a median CSM velocity of $v_{\rm CSM}\simeq 1200{~{}\rm km}{~{}\rm s}^{-1}$, the MED time from the merger event to the explosion is $\simeq 6\times 10^{15}{~{}\rm cm}/v_{\rm CSM}\simeq 5\times 10^{7}{~{}\rm s}$, and so I estimate $t_{\rm MED}\simeq 1-2{~{}\rm yr}$.
During the merger event, the binary system ejects the CSM. The escape velocity from the merger system, $\approx 5000{~{}\rm km}{~{}\rm s}^{-1}$, as Tsalapatas et al. (2025) note, is much larger than the CSM expansion velocity $v_{\rm CSM}\simeq 1200{~{}\rm km}{~{}\rm s}^{-1}$. I suggest that most of the CSM mass was lost in and near the equatorial plane through the second Lagrange point, i.e., on the side of the less massive WD. Mass lost from the second Lagrange point in strong binary interaction might be at a terminal velocity much slower than the escape velocity from the system, or even stays bound (e.g, Hubová & Pejcha 2019 and references therein).
@@ 3 The problems with some other scenarios
The CD scenario, which incorporates a MED as an essential ingredient, faces two challenges in explaining the properties of SN 2020aeuh; however, these properties do not completely rule it out for SN 2020aeuh. The first is that the common envelope evolution that brings the old WD to merge with the core of an asymptotic giant branch star ejects a hydrogen-rich envelope. The WD-core merger might occur in rare cases later, if the WD survives and the helium envelope expands much later to engulf the WD, causing it to merge with the core. Soker & Bear (2023) suggested this channel of the CD scenario to explain the compact helium-rich CSM of SN 2020eyj (Kool et al. 2023). However, it is unclear if a common envelope evolution can take place without helium presence in the ejected CSM. According to the CD scenario, the ejecta will eventually collide, within approximately a thousand years, with a hydrogen-rich nebula that was once a planetary nebula. According to the CD scenario, the asymptotic giant branch that engulfs the WD should have a zero-age main-sequence mass of $\gtrsim 4M_{\odot}$ to force the WD to merge with the core. Thus, some relatively recent star formation must have taken place in the host galaxy of SN 2020aeuh (in the last $\simeq 10^{8}{~{}\rm yr}$. These two challenges make the DD-MED scenario the most suitable for explaining SN 2020aeuh.
@@ 4 Summary
SN 2020aeuh belongs to the group of somewhat overluminous SNe Ia, specifically 1991T-like SNe Ia, that some take to be super-Chandrasekhar (e.g., Fisher et al. 1999), namely, result from a WD with a mass $>M_{\rm Ch}$. Often SNe Ia-CSM are overluminous (1991T-like SNe Ia; e.g., Phillips et al. 2024), but otherwise are normal SNe Ia. In the lonely WD scenarios, what separates normal SNe Ia that interact with a close CSM from other normal SNe Ia is a relatively short MED time. I raise the possibility that the reason for the short MED time of SNe Ia-CSM and their overluminous brightness is the mass of the exploding lonely WD being on the upper end of the mass distribution of exploding lonely WDs. Namely, they are indeed somewhat super-Chandrasekhar.
@@
The uncertainties are very large, more than an order of magnitude, and the only certain point is that SNe Ia-CSM-CO are very rare; all types of SNe Ia-CSM are rare (e.g., Terwel et al. 2025a, b). Better estimates of the fraction of such events are for the future.
@@ Acknowledgments
I thank the Charles Wolfson Academic Chair at the Technion for the support.

### v2

@@ Abstract
I suggest the double-degenerate (DD) scenario with a merger-to-explosion delay (MED) time (the DD-MED scenario) of about 1-2 years to explain the rare properties of the recently analyzed type Ia supernova (SN Ia) SN 2020aeuh. The rare properties are the SN Ia ejecta interacting with a carbon-oxygen (CO)-rich circumstellar material (CSM) at approximately 50 days post-explosion. In this DD-MED scenario, two massive CO white dwarfs (WDs), masses of $M_{1}\simeq 1.1M_{\odot}$ and $M_{2}\simeq 1M_{\odot}$, merge to leave a rapidly rotating lonely WD of about the Chandrasekhar mass. The merger process ejects $M_{\rm CSM}\simeq 0.7M_{\odot}$ to form a nonspherical CO-rich CSM. At the explosion, there is a lonely WD and a detached hydrogen- and helium-deficient CSM. Studies proposed the other lonely WD scenario, the core-degenerate (CD) scenario, to explain several specific SNe Ia and SN Ia remnants. SN 2020aeuh is the first particular SN Ia that is attributed to the DD-MED scenario. Besides being slightly brighter than typical SNe Ia and the CSM interaction, SN 2020aeuh is a normal SN Ia. Therefore, this study strengthens the claim of earlier studies, which are based on other arguments, like the properties of SN Ia remnants, that the lonely WD scenarios, i.e., the DD-MED and CD scenarios, might account for most, if not all, normal SNe Ia. These earlier studies also argue that all SN Ia scenarios, whether lonely WD or not, might contribute to peculiar SNe Ia.
@@ 1 INTRODUCTION
The old classification of type Ia supernovae (SNe Ia) scenarios, into the double-degenerate (DD) and single-degenerate (SD) scenarios, is outdated (despite many papers still mentioning this classification as if it is applicable). There are more scenarios that any modeling of an SN Ia must consider. However, there is no consensus on the classification of these scenarios as reviews show (Maoz et al. 2014; Maeda, & Terada 2016; Hoeflich 2017; Livio & Mazzali 2018; Soker 2018, 2019a, 2024a; Wang 2018; Jha et al. 2019; Ruiz-Lapuente 2019; Ruiter 2020; Aleo et al. 2023; Liu, Röpke, & Han 2023; Vinkó, Szalai, & Könyves-Tóth 2023; Ruiter & Seitenzahl 2025). No scenario is entirely free of difficulties in accounting for observations, with some encountering a few challenges, while others face many challenges. Some observations challenge most (e.g., Pearson et al. 2024) or all (e.g., Wang et al. 2024) scenarios. A challenging observation, for example, is the relation of the $\gamma$-ray escape time to the 56Ni mass (e.g., Schinasi-Lemberg & Kushnir 2025; Sharon & Kushnir 2025; Sharon, Kushnir, & Schinasi-Lemberg 2025). Studies have added more sub-scenarios or branches of the main scenarios, including the HeCO hybrid channel of the DD scenario (e.g., Perets et al. 2019; Zenati, Toonen, & Perets 2019; Zenati et al. 2023 ) common-envelope wind model (Meng & Podsiadlowski 2017; Cui et al. 2022; Wang & Meng 2025), the carbon–oxygen–neon white dwarf (WD) with a main sequence companion channel of the SD scenario (e.g., Guo et al. 2025), the core-merger detonation model (Ablimit 2021), and new mass transfer prescriptions (e.g., Li, Liu, & Wang 2023). Studies have examined most of these scenarios in recent years, with no emerging leading scenario(s) in the consensus (e.g., Bora et al. 2024; Boos, Townsley, & Shen 2024; Bregman et al. 2024; Casabona & Fisher 2024; DerKacy et al. 2024; Joshi, Strolger, & Zenati 2024; Ko et al. 2024; Kobashi et al. 2024; Lim et al. 2024; Mehta et al. 2024; Palicio et al. 2024; Phillips et al. 2024; Soker 2024b; Uchida et al. 2024; Burmester et al. 2025; Court et al. 2025; Gaba et al. 2025; Glanz et al. 2025; Griffith et al. 2025; Hoogendam et al. 2025a, b; Ito et al. 2025; Iwata & Maeda 2025; Kumar, Prust, & Bildsten 2025; Magee et al. 2025; Michaelis & Perets 2025; O’Hora et al. 2025; Pan, Ruiz-Lapuente, & González Hernández 2025; Pollin et al. 2025; Simotas, Bildsten, & Prust 2025; Wang, Chen, & Pan 2025, for a limited list of papers since 2024).
@@ 2 The DD-MED scenario for SN 2020aeuh
A property that Tsalapatas et al. (2025) deduce from their interaction model and not directly from observables.
@@
Tsalapatas et al. (2025) find the full width at half-maximum intensity of two narrow (relative to the ejecta) spectral lines to be $\simeq 750{~\rm km}{~\rm s}^{-1}$ and $\simeq 1250{~\rm km}{~\rm s}^{-1}$, and a third asymmetrical line shows a velocity of $\simeq 1350{~\rm km}{~\rm s}^{-1}$. I, therefore, scale the CSM velocity with $v_{\rm CSM}\simeq 1200{~\rm km}{~\rm s}^{-1}$. Tsalapatas et al. (2025) take the ejecta velocity to be $v_{\rm ej}\simeq 1.2\times 10^{4}{~\rm km}{~\rm s}^{-1}$, and from that estimate the CSM shell to be from $R_{\rm CSM,inner}\simeq 3\times 10^{15}{~\rm cm}$ to $R_{\rm CSM,outer}\simeq 9\times 10^{15}{~\rm cm}$. The CSM velocity is about an order of magnitude slower. From the explosion until the ejecta catches up with the CSM, the radius of the CSM increases by a factor of $\simeq 1.1$, i.e., by $\simeq 10\%$. Therefore, for the accuracy of this calculation, the exact expansion of the CSM from explosion to the time of ejecta-CSM interaction is of little significance. For a median CSM velocity of $v_{\rm CSM}\simeq 1200{~\rm km}{~\rm s}^{-1}$, the time from CSM ejection to explosion, namely, the MED time, is $t_{\rm MED}\simeq 0.5(R_{\rm CSM,inner}+R_{\rm CSM,outer})/v_{\rm CSM}$. The MED time from the merger event to the explosion is, therefore, $t_{\rm MED}\simeq 6\times 10^{15}{~\rm cm}/1,200{~\rm km}{~\rm s}^{-1}=5\times 10^{7}{~\rm s}$, and so I estimate $t_{\rm MED}\simeq 1-2{~\rm yr}$. The value of $v_{\rm CSM}$ is the average velocity of the CSM. The inner parts of the CSM are slower. For a delay time of $t_{\rm MED}\simeq 1.5{~\rm yr}$, an inner CSM velocity of $v_{\rm CSM,inner}\simeq 650{~\rm km}{~\rm s}^{-1}$ gives the observed inner CSM radius of $R_{\rm CSM,inner,D}\simeq v_{\rm CSM,inner}t_{\rm MED}\simeq 3.1\times 10^{15}{~\rm cm}$. This is the DD-MED explanation for the inner CSM radius.
During the merger event, the binary system ejects the CSM. The escape velocity from the merger system, $\approx 5000{~\rm km}{~\rm s}^{-1}$, as Tsalapatas et al. (2025) note, is much larger than the CSM expansion velocity $v_{\rm CSM}\simeq 1200{~\rm km}{~\rm s}^{-1}$. I suggest that most of the CSM mass was lost in and near the equatorial plane through the second Lagrange point, i.e., on the side of the less massive WD. Mass lost from the second Lagrange point in strong binary interaction might be at a terminal velocity much slower than the escape velocity from the system, or even stays bound (e.g, Hubová & Pejcha 2019 and references therein). Raskin & Kasen (2013) simulated the merger of two WDs in the DD scenario, and found a tidal tail expanding at $\approx 2000{~\rm km}{~\rm s}^{-1}$. They found the ejected mass to be $\approx 0.001-0.005M_{\odot}$. The lonely WD scenario lets the merger product relax and eject more mass.
@@ 3 The problems with some other scenarios
The CD scenario, which incorporates a MED as an essential ingredient, faces two challenges in explaining the properties of SN 2020aeuh; however, these properties do not completely rule it out for SN 2020aeuh. The first is that the common envelope evolution that brings the old WD to merge with the core of an asymptotic giant branch star ejects a hydrogen-rich envelope. The WD-core merger might occur in rare cases later, if the WD survives and the helium envelope expands much later to engulf the WD, causing it to merge with the core. Soker & Bear (2023) suggested this channel of the CD scenario to explain the compact helium-rich CSM of SN 2020eyj (Kool et al. 2023). However, it is unclear if a common envelope evolution can take place without helium presence in the ejected CSM. According to the CD scenario, the ejecta will eventually collide, within approximately a thousand years, with a hydrogen-rich nebula that was once a planetary nebula. According to the CD scenario, the asymptotic giant branch that engulfs the WD should have a zero-age main-sequence mass of $\gtrsim 4M_{\odot}$ to force the WD to merge with the core. Thus, some relatively recent star formation must have taken place in the host galaxy of SN 2020aeuh (in the last $\simeq 10^{8}{~\rm yr}$). These two challenges of the CD scenario make the DD-MED scenario the most suitable for explaining SN 2020aeuh.
@@ 4 Summary
SN 2020aeuh belongs to the group of somewhat overluminous SNe Ia, specifically 1991T-like SNe Ia, that some take to be super-Chandrasekhar (e.g., Fisher et al. 1999), namely, result from a WD with a mass $M_{\rm Ch}$. Often SNe Ia-CSM are overluminous (1991T-like SNe Ia; e.g., Phillips et al. 2024), but otherwise are normal SNe Ia. In the lonely WD scenarios, what separates normal SNe Ia that interact with a close CSM from other normal SNe Ia is a relatively short MED time. I raise the possibility that the reason for the short MED time of SNe Ia-CSM and their overluminous brightness is the mass of the exploding lonely WD being on the upper end of the mass distribution of exploding lonely WDs. Namely, they are indeed somewhat super-Chandrasekhar.
@@
I scale with $\Delta t_{\rm pe}=10{~\rm yr}$, which is the typical time scale that observations are searching for late ejecta-CSM interaction after the SN Ia explosions The uncertainties are very large, more than an order of magnitude, and the only certain point is that SNe Ia-CSM-CO are very rare; all types of SNe Ia-CSM are rare (e.g., Terwel et al. 2025a, b). Better estimates of the fraction of such events are for the future.
@@ Acknowledgments
I thank an anonymous referee for useful comments. I thank the Charles Wolfson Academic Chair at the Technion for the support.